Find sum to n terms of series 3.1² + 4.2² + 5.3²......

Grade:11

1 Answers

Priyanka Bharti
5 years ago
3×1^2 + 4×2^2 + 5×3^2 +.......
= (1+2)×1^2 + (2+2)×2^2 + (3+2)×3^2 + ..    
= 1^3 +2×1^2 + 2^3 + 2×2^2 + 3^3 + 2×3^2 +...  . 
= ( 1^3 + 2^3 + 3^3 + . . . . .) +2(1^2 + 2^2 + 3^2 +... . )
Then apply formula and get the solution